OpenStudy (anonymous):

what's the equation for this? Audrey is 5/6 as old as Bettina. Five years ago, Audrey was 4/5 as old as Bettina was then. What are their present ages?

Directrix (directrix):

Let Audrey's current age be denoted by x. Let Bettina's current age be denoted by y. -------- Audrey is 5/6 as old as Bettina. So, x = (5/6) * Five years ago, Audrey was 4/5 as old as Bettina was then. Five years ago, Audrey was (x -5) years old. Five years ago, Bettina was (y - 5) years old. So, (x - 5) = (4/5) (y - 5) ** x = (5/6) * (x - 5) = (4/5) (y - 5) ** ------------------ Solve these two equations in two unknowns simultaneously. x (Audrey) = 25 and y (Bettina) = 30 @danielle02

OpenStudy (anonymous):

thanks @Directrix

Directrix (directrix):

Glad to help. Any other math problems?